# Integrated population model reveals human and environment driven changes in Baltic ringed seal (Pusa hispida botnica) demography and behavior

### v1

@@ 2.5.3 Predictive simulations
Using posterior samples of the model parameters, we projected the population dynamics of Baltic ringed seals in the Bothnian Bay over the next 15 years under four management scenarios: Scenario 1 maintains current hunting quotas; Scenarios 2 and 3 implement annual increases of 15 and 35 licenses, respectively, for both Finland and Sweden; and Scenario 4 introduces a one-time reduction of 270 licenses in each country, equivalent to a decrease of approximately 75%. These scenarios were applied starting from 2024. For each simulated year, the ice extent in April was randomly sampled from historically observed results. We ran 8,000 simulations of each scenario, corresponding to each posterior sample of the model parameters.
@@ 3 Results
Posterior median estimates from our model indicate that the number of ringed seals in the Bothnian Bay has increased from 4,800 (95%CI: 3,500 - 6,500) in 1988 to 26,700 (20,400 - 35,800) in 2023 (Figure 4a). Driven by an estimated rise in ringed seal birth rates from 0.32 (0.13 - 0.54) to 0.75 (0.68 - 0.81) (Figure 4b), we estimated that the population growth rate increased from 1.6% (-1.7% - 4.4%) in 1988 to a peak of 6.7% (5.6% - 8.1%) in 2015 before declining to as low as 3.9% (2.5% - 5.5%) in 2021 due to the re-introduction of seal hunting (Figure 4c). The dominant eigenvalue of the population projection matrix (Section 2.5.2) suggests that under ideal conditions, the Bothnian Bay population can achieve a maximum growth rate of about 7.0% (6.0% - 8.7%) (Figure 4c), with a corresponding net reproduction rate of 3.3 (2.3 - 4.6) female pups produced per lifetime.
@@
We estimated that pregnancy rates may have been as low as 0.26 (0.02 - 0.54) during the 1970s, and could reach a maximum of 0.81 (0.74 - 0.89) in a healthy Bothnian Bay population. The current pregnancy rate was estimated to be 0.78 (0.72 - 0.84) (Appendix S1: Figure S1), although the birth rate in the spring was 0.75 (0.68 - 0.81). It is therefore possible that about 4% (0% - 8%) of pregnancies are aborted between the fall and spring. The reproductive status of ringed seals is commonly assessed using placental scars and CA. We estimated that by late April, approximately 8% (1% - 15%) of placental scars from births may have already faded sufficiently to become undetectable. In contrast, we estimated that most CA were easily visible between April and June, and were correctly identified 97% (91% - 99%) of the time. In addition to seals that had recently given birth, our estimates indicate that between April and June, CA can also be found in 7% (0% - 24%) of seals that had not given birth.
@@
Our model was not informative on the carrying capacity of the population, indicating that density dependent effects were not detectable at the present population size. Hence, the posterior distribution of the carrying capacity corresponded to our prior assumptions based on historical estimates of the population size (Appendix S2). Assuming density dependence primarily affects fecundity, birth rates may eventually decline to as low as 0.24 (0.17 - 0.35) (Figure 4b). It is worth noting that this latter result is independent of the estimated carrying capacity (Equation 25).
@@
Improvements in reproductive rates have implied a substantial change in the age structure of the population (Figure 4d). Our estimates suggest that the proportion of pups in the population may have increased from 11% (5% - 15%) in 1988 to 18% (16% - 21%) in 2023. During the same period, we estimated a decline in the proportion of adults from 67% (54% - 84%) to 49% (42% - 55%). Assuming density dependence primarily affects fecundity, the age structure at carrying capacity can be expected to consist of 9% (7% - 11%) pups, 18% (12% - 30%) sub-adults and 74% (60% - 81%) adults. If density dependence affects pup survival in addition to fecundity, the proportion of pups at carrying capacity will be higher.
@@
We did not find evidence for a skewed sex ratio which, driven by our prior assumptions, was reflected in similar mortality rates between the sexes in the absence of hunting (Appendix S2). The probability that pups survive all mortality sources other than hunting during their first year was estimated to be 0.62 (0.41 - 0.87), which increased to 0.85 (0.76 - 0.95) at age one and to 0.95 (0.91 - 0.97) by maturity (Figure 5a).
@@
We found significant juvenile bias in bycatch from small-scale fisheries, with pups nearly 8 (4 - 16) times more likely to be bycaught than one year olds (Figure 5b). Male pups were also about 40% (1% - 88%) more likely to be bycaught than female pups. We estimated that Swedish hunting during the spring was representative of population demography (Figure 5c). In contrast, the fall hunt was heavily aimed at adults of both sexes, with a slight bias towards males (Figure 5d). We estimated that pups were the most vulnerable age group to hunting in Finland (Figure 5e). Sub-adults were also generally underrepresented in the Finnish hunting bag.
@@
We estimated that about 60% (48% - 74%) of sub-adult and adult seals and 35% (11% - 60%) of pups may be visible on ice during aerial surveys that coincide with extensive ice cover. The former estimate for sub-adults and adults was influenced by the prior distribution, which had a slightly higher median of 65% (Appendix S2). Historical ice conditions in the Bothnian Bay suggest that approximately 51% (39% - 65%) of the total population may be visible during aerial surveys carried out in typical ice conditions (Figure 6b). We found that as ice cover diminishes, the fraction of hauled-out seals may gradually decline to as low as 44% (30% - 62%). However, this proportion was found to increase dramatically to as much as 91% (71% - 99%) when the ice cover declined below a critical threshold of about 13,000 km2 (Figure 6c).
@@
An analysis of the impact of hunting on population growth revealed that at present, hunting as few as 1,200 seals per year may be sufficient to cause population decline (Figure 7a), although the uncertainty around this estimate was high (1,200 - 2,600). Achieving a growth rate of 7%, the benchmark set by the Baltic Marine Environment Protection Commission (HELCOM) for designating good status, might prove challenging with any level of hunting activity (Figure 7b), especially if more than 340 seals are hunted annually. Posterior predictive simulations suggested that maintaining hunting quotas at their current levels of 350 and 375 seals per year in Sweden and Finland, respectively, may allow for a slight improvement in the population growth rate over the next 10-15 years (Figure 7b). In contrast, increasing quotas by 15 licenses per year in each of Sweden and Finland may cause a slight decline in the growth rate, and an increase of 35 licenses per year in each country may significantly slow down population growth, and may even lead to population decline within the next 15 years.
@@ 4 Discussion
So far, different sources of long-term monitoring data on Baltic ringed seals have been analyzed separately to assess population growth, reproductive health and other relevant metrics of population status. A major limitation of this approach is its inability to leverage all available information, as the synergies between different data sources are not exploited (schaub2011; zipkin2018). This limitation has become more apparent than ever during the last decade, when unexpected trends and large fluctuations in population counts from aerial surveys made it impossible to obtain reliable estimates of population growth from survey data alone. Growth in closed populations, however, is ultimately a consequence of births and deaths. Thus, data related to reproduction and hunting also contain information on the growth rate. By incorporating all available information into a mechanistic IPM, we were able to address recent challenges in ringed seal monitoring in the Bothnian Bay. Our model showed a good fit to all available data, and was able to estimate a large number of ecologically important parameters.
@@
On the other hand, a small number of unusually large estimates obtained in recent years were explained surprisingly well by an increase in the fraction of seals hauling-out during aerial surveys conducted in poor ice conditions, supporting earlier speculations (HELCOMindicator). It is not clear why seals would haul-out in significantly greater numbers under such circumstances. One possibility is that as the ice begins to break up, sub-adult age classes haul-out on the ice to a larger extent and are thus more visible during the surveys (crawford2012; Oksanen2015). As adult seals relinquish territorial behavior typical of winter (Harkonen1992; Harkonen1998; Kelly2010), sub-adults might find opportunities to haul-out in greater numbers. This may additionally result in large congregations of seals, potentially contributing to the high variability in survey estimates (Harkonen1998; Linden2011).
@@
Since the sub-adult age classes account for about a third of the total population, it seems unlikely that an increase in their visibility alone can account for a nearly two-fold increase in population size estimates. The sharp increase in the proportion of visible seals during low ice cover may also be attributed to an increase in the cost of foraging. During the molting period, ringed seals are likely to derive greater benefit from maintaining elevated skin temperatures than from consuming resources (thometz2021). Diminishing sea ice will most likely result in increased distances between optimal haul-out sites and productive foraging areas, and the trade-offs associated with foraging could compel seals to decrease the frequency with which they venture out to sea. Moreover, seals may anticipate that the molting period could be interrupted prematurely, causing them to spend longer periods of time out of the water to ensure that they attain adequate levels of solar exposure before the ice melts completely.
@@
While our findings demonstrate a clear effect of ice cover on aerial survey estimates, establishing haul-out behavior as the causal mechanism demands further investigations, including model comparisons to evaluate alternative hypotheses. Regardless of the mechanism, the surge in survey estimates during low ice cover must be a transient phenomenon, as the number of seals on ice must evidently be zero when no sea ice is present. Extreme inter-annual variability in population counts from aerial surveys may therefore be expected following mild winters (Figure 6c), which could pose additional challenges for future ringed seal monitoring efforts in the Bothnian Bay. As ice conditions in the Baltic Sea continue to deteriorate (meier2004; haapala2015), understanding the factors that affect ringed seal haul-out behavior will become increasingly important for reliable monitoring. Our study demonstrates that aerial surveys can be a valuable source of information to this end, and survey estimates that appear unreliable at first may provide important insights on ringed seal behavior. We therefore emphasize the need for continued monitoring efforts, as well as the need for detailed mechanistic models that go beyond simple analyses of trends.
@@
Statements regarding the total population size of ringed seals require an estimate of the proportion of seals that are hauled-out during the surveys. Haul-out behavior is likely to vary by age (smith1970; lydersen1993; crawford2012; Oksanen2015), but little is known to quantify age-specific differences. Most studies have focused on sub-adult and adult ringed seals, which suggested that 50% to 84% of seals may be visible during surveys (smith1970; Harkonen1992; born2002; Kelly2010). However, pups molt much sooner than older seals, and may spend most of their time in the water by the time they are 5 weeks old (smith1970; lydersen1993). Failing to account for their lower visibility during the surveys may underestimate the total population size. We found that following typical winters, roughly half of the population may be visible on ice during surveys, which corresponds to the lower end of previous estimates, and is similar to what has been proposed by smith1970. The number of ringed seals in the Bothnian Bay may therefore be larger than previously expected (Sundqvist2012).
@@
Little is known about the survival probabilities of ringed seal pups, but previous estimates range from as low as 0.34 in lake Saimaa (kokko1998) to 0.6-0.7 in the Arctic (smith1970; Kelly1988). A survival probability as high as 0.8 has been suggested for unharvested Arctic populations (smith1970). For Baltic ringed seals, a pup survival probability of 0.65 has previously been assumed (Sundqvist2012), which is very similar to our posterior median estimate. However, the uncertainty around our estimate was high. Baltic ringed seal pups are safe from large predators that are common in the Arctic such as polar bears (Ursus maritimus) and killer whales (Orcinus orca), which might lead to higher survival probabilities compared to the Arctic. On the other hand, higher bycatch mortality and greater chance of early snow and ice melt in the Baltic Sea could compensate for the limited presence of predators. Snow and ice conditions in the Baltic Sea are expected to decline in the future, which will most likely lead to increased pup mortality (smith1975; meier2004; Ferguson2005; Kelly2010; Sundqvist2012; Reimer2019).
@@
Although previous estimates of sub-adult survival probabilities are rare, our estimate aligns closely with the 0.81-0.93 range reported in earlier studies (Sundqvist2012; koivuniemi2019). Our posterior median estimate of 0.95 for adult survival probability corresponds to the upper end of previous estimates for ringed seals, which range from 0.88-0.96 (durant1986; Kelly1988; kokko1998; Sundqvist2012; Reimer2019; koivuniemi2019), and is consistent with previous assumptions for Baltic ringed seals (Sundqvist2012). Similarly high adult survival probabilities have also been estimated for Saimaa ringed seals (koivuniemi2019) and British grey seals (Halichoerus grypus) (thomas2019).
@@
We estimated a historical minimum pregnancy rate that was strongly in agreement with some empirical estimates from the late 1970s (helle1980). However, the uncertainty around our estimate was high due to limited data on female reproductive status during the initial years of our study period. Our estimates indicate that the current pregnancy rate may be nearly the same as the maximum attainable by a healthy Bothnian Bay population. Thus, the population appears to have almost fully recovered from the effects of organochlorine contamination. Assessments of the reproductive health of Baltic ringed seals is currently based on a tentative pregnancy rate threshold of 0.9 that was established by HELCOM based on studies of seal populations elsewhere (HELCOMindicator_reproduction). Our results suggest that this threshold may be too high even for a healthy Bothnian Bay population, possibly because the Bothnian Bay is a relatively unproductive environment (Kauhala2019). A threshold between 0.80-0.85 may be a more realistic target. As the population continues to grow, rising intra-specific competition may cause pregnancy rates to decrease. Body conditions of Baltic ringed seals may already be declining due to poor nutritional status (Kauhala2019), although it is unclear whether population density is a contributing factor.
@@
The most reliable assessments of reproductive rates will be based on the presence of embryos in female seals sampled in the fall. Assessments based on placental scars may underestimate the pregnancy rate, as our results indicate that some scars that were present during parturition may have already faded by the end of April. Many of these scars are likely to be from early births in February. Additionally, a small fraction of fetuses may be aborted in early pregnancy, and placental scar from these pregnancies may have also faded by spring. Accuracy of pregnancy rate estimates from placental scars can be improved either by using the earliest available samples in the spring, or by adjusting the estimates to account for the effect of fading. Assessments based on CA may be more difficult to interpret since a CA could also be present in seals that had an infertile estrous cycle (boyd1984). However, we found that CA may yield very accurate estimates of the pregnancy rate when the pregnancy rate is high.
@@
It is possible that the quotas in 2022 were close to a critical threshold beyond which population growth could no longer be maintained. However, this threshold will increase in the near future provided that the population continues to grow (Figure 7a). Current annual hunting quotas are established at 350 and 375 seals in Sweden and Finland, respectively. Our simulation study suggested that while increasing quotas by fewer than 35 licenses per year in each of Sweden and Finland is unlikely to cause population decline within the next 10-15 years, increases in excess of 15 per year in each country may further decrease the population growth rate (Figure 7b). Our model assumed that there is sufficient demand for seal hunting such that hunting effort will increase proportionally to hunting quotas. While this assumption will not hold for large quotas, it remains appropriate within a precautionary framework.
@@
It should be emphasized that, due to the short time frame of available demographic data and the opportunistic nature of sampling, our model did not take into account likely effects of climate change and environmental stochasticity on ringed seal demography, which may lead to higher pup mortality (meier2004; Ferguson2005; Sundqvist2012; Reimer2019). Thus, our simulation results should be viewed as a comparison of the potential impact of different hunting quota decisions, rather than as forecasts of future population dynamics. As climate change is expected to lead to further declines in snow and ice cover in the Baltic Sea (meier2004; haapala2015), understanding its effects on ringed seal demography is essential for making longer term predictions.

### v2

@@ 2.5.4 Predictive simulations
Using posterior samples of the model parameters, we projected the population dynamics of Baltic ringed seals in the Bothnian Bay over the next 15 years. Since our model lacked an explicit link between demographic parameters and sea ice conditions, we chose not to include climate change projections in future simulations to avoid drawing misleading conclusions. Instead, we focused on comparing the effects of alternative hunting quota scenarios. We considered four management scenarios: Scenario 1 maintains current hunting quotas; Scenarios 2 and 3 implement annual increases of 15 and 35 licenses, respectively, for both Finland and Sweden; and Scenario 4 introduces a one-time reduction of 270 licenses in each country, equivalent to a decrease of approximately 75%. These scenarios were applied starting from 2024. For each simulated year, the ice extent in April was randomly sampled from historically observed results. We ran 8,000 simulations of each scenario, corresponding to each posterior sample of the model parameters, using R Statistical Software version 4.2.0 (Rteam2021).
@@ 3 Results
Posterior median estimates from our model indicate that the total number of ringed seals in the Bothnian Bay has increased from 4,700 (95%CI: 3,500 - 6,600) in 1988 to 26,700 (20,200 - 36,300) in 2023 (Figure 4a). Driven by an estimated rise in ringed seal birth rates from 0.32 (0.13 - 0.53) to 0.74 (0.66 - 0.81) (Figure 4b), we estimated that the population growth rate increased from 1.7% (-1.7% - 4.5%) in 1988 to a peak of 6.8% (5.6% - 8.1%) in 2015 before declining to as low as 4.0% (2.5% - 5.5%) in 2021 due to the re-introduction of seal hunting (Figure 4c). The dominant eigenvalue of the population projection matrix suggests that under ideal conditions, the Bothnian Bay population can achieve a maximum growth rate of about 7.0% (5.9% - 8.8%) (Figure 4c), with a corresponding net reproduction rate of 3.3 (2.3 - 4.6) female pups produced per lifetime (Section 2.5.3).
We estimated that pregnancy rates may have been as low as 0.27 (0.02 - 0.56) during the 1970s, and could reach a maximum of 0.83 (0.75 - 0.93) in a healthy Bothnian Bay population. Reproductive rates were primarily informed by direct data on reproductive signs, but the aerial survey data also provided some indication of an increasing trend in reproduction over time (Supplement 1: Figure S7). The current pregnancy rate was estimated to be 0.82 (0.74 - 0.89) (Supplement 1: Figure S2), although the birth rate in the spring was 0.74 (0.66 - 0.81). It is therefore possible that about 9% (1% - 21%) of pregnancies are aborted between the fall and spring. However, the pregnancy estimates were somewhat sensitive to assumptions about the importance of late-term abortions. When late-term abortions were assumed to be relatively rare, corresponding to the model with decreasing rate of pregnancy losses (Supplement 4), the estimated current pregnancy rate was 0.78 (0.72 – 0.84), with 4% (0 – 8%) of pregnancies failing between fall and winter. The choice of the model for the pregnancy losses was insignificant for other results.
@@
The reproductive status of ringed seals is commonly assessed using placental scars and CA. We estimated that by late April, approximately 7% (1% - 15%) of placental scars from births may have already faded sufficiently to become undetectable. In contrast, we estimated that most CA were easily visible between April and June, and were correctly identified 97% (91% - 100%) of the time. In addition to seals that had recently given birth, our estimates indicate that between April and June, CA can also be found in 9% (0% - 26%) of seals that had not given birth.
@@
Our model was not informative on the carrying capacity of the population, indicating that density dependent effects were not detectable at the present population size. Hence, the posterior distribution of the carrying capacity corresponded to our prior assumptions based on historical estimates of the population size (Supplement 2). Assuming density dependence primarily affects fecundity, birth rates may eventually decline to as low as 0.24 (0.17 - 0.34) (Figure 4b). It is worth noting that this latter result is independent of the estimated carrying capacity (caswell2001).
@@
Improvements in reproductive rates have implied a substantial change in the age structure of the population (Figure 4d). Our estimates suggest that the proportion of pups in the population may have increased from 11% (5% - 15%) in 1988 to 18% (16% - 21%) in 2023. During the same period, we estimated a decline in the proportion of adults from 67% (54% - 84%) to 50% (43% - 55%). Assuming density dependence primarily affects fecundity, the age structure at carrying capacity can be expected to consist of 9% (6% - 11%) pups, 18% (12% - 29%) sub-adults and 74% (61% - 82%) adults.
@@
We did not find evidence for a skewed sex ratio which, driven by our prior assumptions, was reflected in similar mortality rates between the sexes in the absence of hunting (Supplement 2). The probability that pups survive all mortality sources other than hunting during their first year was estimated to be 0.63 (0.42 - 0.87), which increased to 0.86 (0.77 - 0.95) at age one and to 0.95 (0.91 - 0.97) by maturity (Figure 5a). However, the large negative posterior correlation between juvenile and adult survival probabilities suggested that the model had limited capacity to estimate age-specific mortality rates (Supplement 1: Figure S11). Refitting the model with different subsets of the data confirmed that survival probabilities were primarily informed by pre-2012 aerial survey and the reproduction data, both of which lack information on age-specific mortality (Supplement 1: Figure S5).
@@
We found significant juvenile bias in bycatch from small-scale fisheries, with pups nearly 8 (4 - 16) times more likely to be bycaught than one year olds (Figure 5b). Male pups were also about 40% (1% - 88%) more likely to be bycaught than female pups. We estimated that Swedish hunting during the spring was representative of population demography (Figure 5c). In contrast, the fall hunt was heavily aimed at adults of both sexes, with a slight bias towards males (Figure 5d). We estimated that pups were the most vulnerable age group to hunting in Finland (Figure 5e). Sub-adults were also generally underrepresented in the Finnish harvest.
@@
We estimated that about 60% (48% - 74%) of sub-adult and adult seals and 35% (12% - 60%) of pups may be visible on ice during aerial surveys that coincide with extensive ice cover. Historical ice conditions in the Bothnian Bay suggest that approximately 51% (39% - 65%) of the total population may be visible during aerial surveys carried out in typical ice conditions (Figure 6b). We found that as ice cover diminishes, the fraction of hauled-out seals may gradually decline to as low as 44% (30% - 62%). However, this proportion was found to increase dramatically to as much as 91% (70% - 99%) when the ice cover declined below a critical threshold of about 13,000 km2 (Figure 6). Although our mechanistic haul out probability model was relatively complex, all parameters appeared to be well informed, largely by the aerial survey data. The exception was the baseline haul out probability $\hat{\omega}_{s,a}$, which was given an informative prior (Supplement 2) and was only weakly informed by data. Nonetheless, posterior estimates were generally robust to prior choice. Only estimates of total population size, and consequently population growth rate under harvesting, showed moderate sensitivity, with estimates varying by less than 10% under sensible prior choices (Supplement 1: Table LABEL:table:post_post_overlap).
@@
An analysis of the impact of hunting on population growth revealed that at present, hunting as few as 1,200 seals per year may be sufficient to cause population decline (Figure 7a), although the uncertainty around this estimate was high (1,200 - 2,600). Achieving a growth rate of 7%, the benchmark set by the Baltic Marine Environment Protection Commission (HELCOM) for designating good status, might prove challenging with any level of hunting activity (Figure 7b), especially if more than 380 seals are hunted annually. Posterior predictive simulations suggested that maintaining hunting quotas at their current levels of 350 and 375 seals per year in Sweden and Finland, respectively, may allow for a slight improvement in the population growth rate over the next 10-15 years (Figure 7b). In contrast, increasing quotas by 15 licenses per year in each of Sweden and Finland may cause a slight decline in the growth rate, and an increase of 35 licenses per year in each country may significantly slow down population growth, and may even lead to population decline within the next 15 years.
@@ 4 Discussion
So far, different sources of long-term monitoring data on Baltic ringed seals have been analyzed separately to assess population growth, reproductive health and other relevant metrics of population status. A major limitation of this approach is its inability to leverage all available information, as the synergies between different data sources are not exploited (schaub2011; zipkin2018). This limitation has become more apparent than ever during the last decade, when unexpected trends and large fluctuations in population counts from aerial surveys made it impossible to obtain reliable estimates of population growth from survey data alone. Growth in closed populations, however, is ultimately a consequence of births and deaths. Thus, data related to reproduction and hunting also contain information on the growth rate. By incorporating all available information into a mechanistic IPM, we were able to address recent challenges in ringed seal monitoring in the Bothnian Bay. Our model showed a good fit to all available data (Supplement 1), and was able to estimate a large number of ecologically important parameters. Moreover, fitting our model in Stan (stan2018), while requiring some approximations, proved highly efficient, achieving convergence in under an hour on a personal laptop computer despite the large number of parameters.
@@
On the other hand, a small number of unusually large population estimates obtained in recent years were explained surprisingly well by an increase in the fraction of seals hauling out during aerial surveys conducted in poor ice conditions, supporting earlier speculations (HELCOMindicator). It is not clear why seals would haul out in significantly greater numbers under such circumstances. Sea ice conditions during the surveys likely correlate with air temperature, wind speed, and snow depth, which are known to affect ringed seal haul out probability (Harkonen1992; hamilton2018). Snow depth may additionally influence the proportion of seals that are hiding in lairs (Kelly2010). However, as surveys are conducted in late April, when most lairs have collapsed and seals are at the height of molting, and only on clear, calm days, the impact of these variables on population size estimates is likely limited. Another possibility is that the break up of sea ice causes adult seals to relinquish territorial behavior typical of winter, allowing younger sub-adults to haul out in greater numbers (Harkonen1992; Harkonen1998; Kelly2010; Oksanen2015). This may additionally result in large congregations of seals, potentially contributing to the high variability in survey estimates (Harkonen1998; Linden2011).
@@
Since the sub-adult age classes account for about a third of the total population, it seems unlikely that an increase in their visibility alone can account for a nearly two-fold increase in population size estimates. Likewise, immigration from southern sub-populations, which are collectively believed to be nearly an order of magnitude smaller than the Bothnian Bay population (Sundqvist2012), also appears insufficient to explain the observed increase in counts. The increase in the proportion of visible seals during low ice cover could, in part, be attributed to an increase in the cost of foraging. During the molting period, ringed seals are likely to derive greater benefit from maintaining elevated skin temperatures than from consuming resources (thometz2021). Diminishing sea ice will most likely result in increased distances between optimal haul-out sites and productive foraging areas, and the trade-offs associated with foraging could compel seals to decrease the frequency with which they venture out to sea. Moreover, seals may anticipate that the molting period could be interrupted prematurely, causing them to spend longer periods of time out of the water to ensure that they attain adequate levels of solar exposure before the ice melts completely.
@@
While our findings demonstrate a clear effect of ice cover on aerial survey estimates, this conclusion was largely driven by two outliers that corresponded to unusually mild winters. Establishing haul-out behavior as the causal mechanism demands further investigations, including additional data and model comparisons to evaluate alternative hypotheses. Regardless of the mechanism, the number of seals on ice must evidently be zero when no sea ice is present. Extreme inter-annual variability in population counts from aerial surveys may therefore be expected following mild winters (Figure 6c), which could pose additional challenges for future ringed seal monitoring efforts in the Bothnian Bay. As ice conditions in the Baltic Sea continue to deteriorate (meier2004; haapala2015), understanding the factors that affect ringed seal haul-out behavior will become increasingly important for reliable monitoring. Our study demonstrates that aerial surveys can be a valuable source of information to this end, and survey estimates that appear unreliable at first may provide important insights on ringed seal behavior. We therefore emphasize the need for continued monitoring efforts, as well as the need for detailed mechanistic models that go beyond simple analyses of trends. Future IPMs could also incorporate individual level data collected from telemetry devices with dry/wet sensors, enabling direct estimation of haul out probabilities (london2024).
@@
Statements regarding the total population size of ringed seals require an estimate of the proportion of seals that are hauled out during the surveys. Haul-out behavior is likely to vary by age (smith1970; lydersen1993; crawford2012; Oksanen2015), but little is known to quantify age-specific differences. Most studies have focused on sub-adult and adult ringed seals, which suggested that 50% to 84% of seals may be visible during surveys (smith1970; Harkonen1992; born2002; Kelly2010). However, pups undergo their first molt and shed their lanugo hair before the molting season of older seals. Thus they spend most of their time in the water by the time they are 5 weeks old (smith1970; lydersen1993). Failing to account for their lower visibility during the surveys may lead to an underestimation of the total population size. We found that following typical winters, roughly half of the population may be visible on ice during surveys, which corresponds to the lower end of previous estimates, and is similar to what has been proposed by smith1970. The number of ringed seals in the Bothnian Bay may therefore be larger than previously expected (Sundqvist2012). However, our results were influenced by prior assumptions about haul out probabilities in abundant ice conditions. Detailed studies on the molting behavior of Baltic ringed seals will be crucial for obtaining reliable estimates of total abundance.
@@
Little is known about the survival probabilities of ringed seal pups, but previous estimates range from as low as 0.34 in lake Saimaa before active conservation efforts, (kokko1998) to 0.6-0.7 in the Arctic (smith1970; Kelly1988). A survival probability as high as 0.8 has been suggested for unharvested Arctic populations (smith1970). For Baltic ringed seals, a pup survival probability of 0.65 has previously been assumed (Sundqvist2012), which is very similar to our posterior median estimate. However, the uncertainty around our estimate was high due to weak parameter identifiability (Supplement 1: Figure S11). Baltic ringed seal pups are safe from large predators that are common in the Arctic such as polar bears (Ursus maritimus) and Arctic foxes (Vulpes lagopus), which might lead to comparably higher survival probabilities (stirling2004). On the other hand, higher bycatch mortality and greater chance of early snow and ice melt in the Baltic Sea could compensate for the limited presence of predators. Snow and ice conditions in the Baltic Sea are expected to decline in the future, which will most likely lead to increased pup mortality (smith1975; meier2004; Ferguson2005; Kelly2010; Sundqvist2012; Reimer2019).
@@
Although previous estimates of sub-adult survival probabilities are rare, our estimate aligns closely with the 0.81-0.93 range reported in earlier studies (Sundqvist2012; koivuniemi2019). Our posterior median estimate of 0.95 for adult survival probability corresponds to the upper end of previous estimates for ringed seals, which range from 0.88-0.96 (durant1986; Kelly1988; kokko1998; Sundqvist2012; koivuniemi2019; Reimer2019), and is consistent with previous assumptions for Baltic ringed seals (Sundqvist2012). Similarly high adult survival probabilities have also been estimated for Saimaa ringed seals (koivuniemi2019) and British grey seals (Halichoerus grypus) (thomas2019). However, the high posterior correlation between juvenile and adult survival probabilities suggests that these parameters were not individually identifiable (Supplement 1: Figure S11). Direct data on survival, such as from mark-recapture studies, will be needed to obtain precise age-specific estimates. Such data can additionally support the estimation of overall abundance and, in turn, haul out probabilities.
@@
We estimated a historical minimum pregnancy rate that was strongly in agreement with some empirical estimates from the late 1970s (helle1980). However, the uncertainty around our estimate was high due to limited data on female reproductive status during the initial years of our study period. Our estimates indicate that the current pregnancy rate may be nearly the same as the maximum attainable by a healthy Bothnian Bay population. Thus, the population appears to have almost fully recovered from the effects of organochlorine contamination. Assessments of the reproductive health of Baltic ringed seals is currently based on a tentative pregnancy rate threshold of 0.9 that was established by HELCOM based on studies of seal populations elsewhere (HELCOMindicator_reproduction). Our results suggest that this threshold may be too high even for a healthy Bothnian Bay population, possibly because the Bothnian Bay is a relatively unproductive environment (Kauhala2019). A threshold between 0.80-0.85 may be a more realistic target. As the population continues to grow, rising intra-specific competition may cause pregnancy rates to decrease. Body conditions of Baltic ringed seals may already be declining due to poor nutritional status (Kauhala2019), although it is unclear whether population density is a contributing factor. Our model did not provide information on density-dependent effects on reproductive rates, either because the current population size is substantially below carrying capacity, or because other demographic parameters, such as pup survival, were more strongly affected. Long-term predictions of ringed seal population dynamics will require detailed consideration of density-dependent mechanisms, which would greatly benefit from additional demographic data such as long-term mark-recapture studies (e.g. hiby2007; koivuniemi2019). Understanding seasonal patterns of reproductive failure and their links to population density and nutritional stress would also be valuable for both assessing current reproductive health and improving future population projections.
@@
The most reliable assessments of reproductive rates will be based on the presence of embryos in female seals sampled in the fall. Unfortunately, the current sample size for this measure remains relatively small, with substantial inter-annual variability in the observed pregnancy rate. Assessments based on placental scars may underestimate the pregnancy rate, as our results indicate that some scars that were present during parturition may have already faded by the end of April. Many of these scars are likely to be from early births in February. Additionally, a small fraction of fetuses may be aborted in early pregnancy, and placental scar from these pregnancies may have also faded by spring. Accuracy of pregnancy rate estimates from placental scars can be improved either by using the earliest available samples in the spring, or by adjusting the estimates to account for the effect of fading. Assessments based only on CA may be more difficult to interpret since a CA could also be present in seals that had an infertile estrous cycle (boyd1984). However, we found that CA may yield very accurate estimates of the pregnancy rate when the pregnancy rate is high. An additional challenge when estimating pregnancy rates is accounting for potential sampling biases. Most assessments, including the present study, assume that reproductive signs in harvested or bycaught females is representative of the entire female population. This assumption may introduce bias if reproductive status affects a female’s likelihood of being harvested or bycaught. For instance, preferential harvesting of pups can inflate reproductive rate estimates if their mothers are harvested with them. Similarly, if postpartum females forage differently due to reduced energy reserves or molt at different locations, their susceptibility to hunting or bycatch may differ. Detailed studies on seal and hunter behavior will be essential for addressing potential sampling biases.
@@
It is possible that the quotas in 2022 were close to a critical threshold beyond which population growth could no longer be maintained. However, this threshold will increase in the near future provided that the population continues to grow (Figure 7a). Current annual hunting quotas are established at 350 and 375 seals in Sweden and Finland, respectively. Our simulation study suggested that while increasing quotas by fewer than 35 licenses per year in each of Sweden and Finland is unlikely to cause population decline within the next 10-15 years, increases in excess of 15 per year in each country may further decrease the population growth rate (Figure 7b). Our model assumed that there is sufficient demand for seal hunting such that harvest rates will increase proportionally to hunting quotas. While this assumption will not hold for large quotas, it remains appropriate within a precautionary framework.
@@
It should be emphasized that, due to the short time frame of available demographic data and the opportunistic nature of sampling, our model did not take into account likely effects of climate change and other environmental variables, such as prey availability, on demographic parameters like pup mortality and reproductive rates (meier2004; Ferguson2005; Sundqvist2012; Reimer2019). Thus, our simulation results should be viewed as a comparison of the potential impact of different hunting quota decisions, rather than as forecasts of future population dynamics. As climate change is expected to lead to further declines in snow and ice cover in the Baltic Sea (meier2004; haapala2015), understanding its effects on ringed seal demography is essential for making longer term predictions. Future IPMs could explicitly link ringed seal survival and reproduction to environmental variables such as snow and ice conditions to predict population responses to various climate change scenarios (e.g. Sundqvist2012; Reimer2019). However, estimating parameters of such models will likely require direct data on survival probabilities, such as from mark-recapture studies.